40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 10 03:44:51 2015, Jenne, Update, LSC, Some small progress, may have DAC problem? PRMI_sb_powers_9Apr2015.png
    Reply  Fri Apr 10 12:09:19 2015, ericq, Update, LSC, Some small progress, may have DAC problem? 
    Reply  Wed Apr 15 21:00:56 2015, Jenne, Update, LSC, DAC fine 
Message ID: 11212     Entry time: Fri Apr 10 03:44:51 2015     Reply to this: 11213   11222
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Some small progress, may have DAC problem? 

Small steps tonight, but all in the forward direction.

On one of my better locks, I saw a kind of weird phenomenon with the PRMI sideband powers versus the carrier powers:

For the last 100 seconds of this plot, I'm all 1f.  Alignment is being handled mostly by Q's DC coupled ITM oplevs, and the transmission QPD ASC loops, although I was trying to adjust the offsets in the ASC loops to improve transmission for a bit.

At the very end, the last 10 seconds or so, the POP110 power goes down, and sits at about half it's maximum value.  POP22 isn't quite as bad, in that it still touches the max, but the RIN is about 50%.  The carrier DC signals (TRX, TRY, POPDC) don't see this huge jump.  I don't think I was touching anything the last few tens of seconds.  I'm not sure yet how I can so significantly lose sideband power, without losing a similar amount of carrier power. 

The ring-ups at about -70sec in the CARM and DARM outs are the bounce mode. 

I tried looking at 2D histograms of different combinations of channels, for the time around -30 seconds where things looked pretty clean.  It looks like the offsets that Q put in last night (+1 for MICH_B and -3 for PRCL_B) are still about right.  The PRCL_IN1 and MICH_IN1 were centered around zero at the maximum power points.  CARM and DARM had small offsets, which I put into the DARM_B and CARM_B filter banks (0.0066 for DARM_B, and 0.027 for CARM_B), although these are small enough that I don't know that they really do anything.


As a break from locking for a little while, I tried to see if I could get the TT3 and TT4 DAC channels to work for me.  I had hoped it would be a quick and easy task, but I'm not seeing signal out.  Since it wasn't working, I decided to go back to locking for the night, and look into the DAC in the daytime.  I want to use one channel as the IN2 input of the CM board, and another as the external modulation input to the Marconi for transfer functions, so I need them to work. 

As a side note on the input to the Marconi situation, it occurred to me that instead of laying a new cable, I can borrow the POP55 heliax.  We don't have a POP55 diode right now, and the other end comes out across the hall from the Marconi, so it would be pretty easy to have a medium-length cable go from ITMX table to the Marconi.  Objections to this? 

Attachment 1: PRMI_sb_powers_9Apr2015.png  448 kB  | Hide | Hide all
PRMI_sb_powers_9Apr2015.png
ELOG V3.1.3-